A man invested $2000, part at 8% and part at 6%.He earned $144 from his investments. How much did he invest at 8%?
Answer by stanbon(75887)

A man invested $2000, part at 8% and part at 6%.He earned $144 from his investments. How much did he invest at 8%?
-----
Quantity Equation: e + s = 2000
Interest Equation:0.08e + 0.06s = 144
-------------------------
Multiply thru the Quantity Equation by 6
Multiply thru the Interest Equation by 100
-------------------------
6e + 6s = 6*2000
8e + 6s = 100*144
-------------------------
Subtract and solve for "e":
2e = 2400
e = $1200 (amt. invested at 8%)
==============
Cheers,
Stan H.
==============